OpenStudy (anonymous):

7m+n =46 m-3n=16 solve using the substitution method

OpenStudy (sid1729):

Substitution means substituting the value of one variable (in terms of the other variable) In this example, let's rewrite equation 2 as: m = 3n + 16. Thus, we can write m in terms of n, and substitute it in equation 1. So equation 1 becomes : 7(3n+16) + n = 46. Now, this equation has only one variable, n, so you can find n from it. Once you find n, put it in the previous equation to find m.

OpenStudy (anonymous):

so would it be 21n+112+n=46 22n+112=46 22n+112-112=46-112 22n=-66 n=-3 m=3(-3)+16 m=-9+16 m=7

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @87corb05 If you followed what @sid1729 said, you would get the answers which you arrived to. Good job, your answer is correct! :)